OpenStudy (anonymous):

I need help with a table of values for f(x)=2x^4 I have a graph. But I don't how how to describe the end behavior.

OpenStudy (aum):

If you need to create a table, substitute a few values for x and calculate the corresponding f(x) value. For example, when x = 2, f(2) = 2 * 2^4 = 2 * 16 = 32. You can try x = 0, 1, 2, 3, etc.

OpenStudy (aum):

Also note that x is raised to an EVEN power. That means whether we put x = +2 or x = -2 we will get the same y-value. So whatever y-values you calculate for the positive x-values will be the same for the corresponding negative x-values.